3+5(2x-4)=-7(5x+2)
Answer by checkley77(12569)

3+5(2x-4)=-7(5x+2)
3+10x-20=-35x-14
10x+35x=-14-3+20
45x=3
x=3/45
x=1/15 answer.
Proof:
3+5(2*1/15-4)=-7(5*1/15+2)
3+5(2/15-4)=-7(5/15+2)
3+10/15-20=-35/15-14
10/15+35/15=17-14
45/15=3
3=3